# On the Impact of Dataset Properties on Membership Privacy of Deep Learning

### v2

@@ Abstract
We apply a state-of-the-art membership inference attack (MIA) to systematically test the practical privacy vulnerability of fine-tuning large image classification models. We focus on understanding the properties of data sets and samples that make them vulnerable to membership inference. In terms of data set properties, we find a strong power law dependence between the number of examples per class in the data and the MIA vulnerability, as measured by true positive rate of the attack at a low false positive rate. We train a linear model to predict true positive rate based on data set properties and observe good fit for MIA vulnerability on unseen data. To analyse the phenomenon theoretically, we reproduce the result on a simplified model of membership inference that behaves similarly to our experimental data. We prove that in this model, the logarithm of the difference of true and false positive rates depends linearly on the logarithm of the number of examples per class. For an individual sample, the gradient norm is predictive of its vulnerability.
@@ 1 Introduction
In this paper, our aim is to apply a state-of-the-art MIA systematically, to help understand practical privacy risks when training deep-learning-based classifiers without DP protections. Our case study focuses on understanding and quantifying factors that influence the vulnerability of non-DP deep learning models to MIA. Since the earliest MIA methods, there has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017). Similarly, there is evidence that models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2023). Based on extensive experimental data over many data sets with varying sizes, we uncover a power law describing the vulnerability remarkably well.
Previously, Tobaben et al. (2023) and Yu et al. (2023) have presented limited studies of how data set properties affect the MIA vulnerability. Tobaben et al. (2023) reported how the MIA vulnerability of few-shot image classification is affected by the number of shots (i.e. the number of examples available per class). Yu et al. (2023) studied how the MIA vulnerability of a class in an image classifier trained from scratch depends on the average individual privacy parameters of that class. Our work significantly expands on both of these by more systematic study and modelling of data set properties and by providing a more detailed look at individual samples, not just classes.
List of contributions In this work we perform an extensive empirical study on the MIA vulnerability of deep learning models. We focus on a transfer learning setting for image classification tasks, where a large pre-trained neural network is fine-tuned on a sensitive data set, and the MIA is performed on the final classifier. We find that the MIA vulnerability has a strong correlation with certain data set properties (Section 3). In Section 4 we analyse the reasons for these findings more thoroughly and we extend the analysis to individual samples in Section 5. The primary contributions of this work are:
@@
    Few-shot MIA: We conduct a comprehensive study of MIA vulnerability (tpr at fixed low fpr) of target models trained using many different data sets with varying sizes and uncover a power law between the number of examples per class and the vulnerability to MIA (see Figure 1).
@@
    Regression model: We utilise our empirical observations to train a regression model to predict MIA vulnerability ($\log(\textsc{tpr}-\textsc{fpr})$ at fixed low fpr) based on examples per class ($\log S$) and number of classes ($\log C$). We show both very good fit on the training data as well as good prediction quality on unseen data from a different feature extractor and when fine-tuning other parameterisations (see Figure 3).
@@
    Theoretical analysis using a simplified model of MIA: To analyse the phenomenon theoretically, we reproduce the result on a simplified model of membership inference that behaves similarly to our experimental data. In this model we can formally prove that $\log(\textsc{tpr}-\textsc{fpr})$ depends linearly on $\log S$ (see Section 4).
@@
    Individual sample MIA: We extend the study to individual samples and find: 1 The gradient norm is predictive of vulnerability of an individual sample. 2 Gradient norms at different stages of learning have different impact, but how depends on the data set (see Section 5).
2 Background and Methods
@@
Membership inference attacks (MIAs) aim to infer whether a particular sample was part of the training set of the targeted model Shokri et al. (2017). Thus, they can be used to determine lower bounds on the privacy leakage of models to complement the theoretical upper bounds obtained through differential privacy.
@@
Likelihood Ratio attack (LiRA; Carlini et al., 2022) While many different MIAs have been proposed Hu et al. (2022), in this work we consider the SOTA: the Likelihood Ratio Attack (LiRA). LiRA assumes an attacker that has black-box access to the attacked model, knows the training data distribution, the training set size, the model architecture, hyperparameters and training algorithm. Based on this information, the attacker can train so-called shadow models (Shokri et al., 2017) which imitate the model under attack but for which the attacker knows the training data set.
@@
LiRA exploits the observation that the loss function value used to train a model is often lower for examples that were part of the training set compared to those that were not. For a target sample $x$, LiRA trains the shadow models: 1 with $x$ as a part of the training set ($x\in\mathcal{D}$) 2 and without $x$ in the training set ($x\notin\mathcal{D}$) . After training the shadow models, $x$ is passed through the shadow models, and based on the losses (or predictions) two Gaussian distributions are formed: one for the losses of $x\in\mathcal{D}$ shadow models, and one for the $x\notin\mathcal{D}$. Finally, the attacker computes the loss for the point $x$ using the model under attack and determines using a likelihood ratio test on the distributions built from the shadow models whether it is more likely that $x\in\mathcal{D}$ or $x\notin\mathcal{D}$. Carlini et al. (2022) proposed an optimisation for performing LiRA for multiple models and points without training a computationally infeasible number of shadow models. We utilise this optimisation.

### v3

# Impact of Dataset Properties on Membership Inference Vulnerability of Deep Transfer Learning

@@ Abstract
We analyse the relationship between privacy vulnerability and dataset properties, such as examples per class and number of classes, when applying two state-of-the-art membership inference attacks (MIAs) to fine-tuned neural networks. We derive per-example MIA vulnerability in terms of score distributions and statistics computed from shadow models. We introduce a simplified model of membership inference and prove that in this model, the logarithm of the difference of true and false positive rates depends linearly on the logarithm of the number of examples per class. We complement the theoretical analysis with empirical analysis by systematically testing the practical privacy vulnerability of fine-tuning large image classification models and obtain the previously derived power law dependence between the number of examples per class in the data and the MIA vulnerability, as measured by true positive rate of the attack at a low false positive rate. Finally, we fit a parametric model of the previously derived form to predict true positive rate based on dataset properties and observe good fit for MIA vulnerability on unseen fine-tuning scenarios.
@@ 1 Introduction
In this paper, our aim is to theoretically understand and apply two state-of-the-art MIAs systematically, to help understand practical privacy risks when fine-tuning deep-learning-based classifiers without DP protections. We focus on transfer learning using fine-tuning because this is increasingly used for all practical applications of deep learning. Our case study focuses on understanding and quantifying factors that influence the vulnerability of non-DP deep transfer learning models to MIA. Since the earliest MIA methods, there has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017). Similarly, there is evidence that models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2023). Feldman & Zhang (2020) showed that neural networks trained from scratch can memorise a large fraction of their training data, while the memorisation is greatly reduced for fine-tuning.
In this paper, we theoretically study the relationship between examples per class, which we denote as shots ($S$), and MIA vulnerability (true positive rate tpr at fixed false positive rate fpr) for a simplified model of fine-tuning and derive a power-law relationship in the form
$\log(\textsc{tpr}-\textsc{fpr})=-\beta_{S}\log(S)-\beta_{0}.$ (1)
We complement the theoretical analysis with extensive experiments over many datasets with varying sizes in the transfer learning setting for image classification tasks and observe the same power law.
Previously, Tobaben et al. (2023) and Yu et al. (2023) have presented limited studies of how dataset properties affect the MIA vulnerability. Tobaben et al. (2023) reported how the MIA vulnerability of few-shot image classification is affected by the number of shots (i.e. the number of examples available per class). Yu et al. (2023) studied how the MIA vulnerability of a class in an image classifier trained from scratch depends on the average individual privacy parameters of that class. Our work significantly expands on both of these by a more systematic study and modelling of dataset properties.
List of contributions We analyze the MIA vulnerability of deep transfer learning using two state-of-the-art MIAs LiRA (Carlini et al., 2022) and RMIA (Zarifzadeh et al., 2024). Both MIAs are based on the assumption that the attacker can train so-called shadow models (Shokri et al., 2017) which imitate the model under attack. We first analyze both MIAs theoretically by introducing a simplified model of membership inference (Section 3). We support our theoretical findings by an extensive empirical study on the MIA vulnerability of deep learning models. We focus on a transfer learning setting for image classification tasks, where a large pre-trained neural network is fine-tuned on a sensitive dataset, and the MIA is performed on the final classifier. We find that the MIA vulnerability has a strong correlation with certain dataset properties (Section 4).
@@
    Per-example vulnerability: We derive per-example LiRA and RMIA vulnerability (tpr at fixed fpr) in terms of MIA attack score distributions and statistics computed from shadow models (see Section 3.3).
@@
    Power-law in simplified model of MI: We formulate a simplified model of membership inference to quantitatively relate dataset properties and MIA vulnerability. For this model, we prove a power-law relationship between the per-example LiRA and RMIA vulnerability and the number of examples per class. We then extend the the power-law relationship to the average-case LiRA and RMIA vulnerability (See Section 3.4).
@@
    Few-shot MIA: We conduct a comprehensive study of MIA vulnerability (tpr at fixed low fpr) in the transfer learning setting for image classification tasks with target models trained using many different datasets with varying sizes and confirm the theoretical power law between the number of examples per class and the vulnerability to MIA (see Figure 1).
@@
    Regression model: We utilise our empirical observations to fit a regression model to predict MIA vulnerability ($\log(\textsc{tpr}-\textsc{fpr})$ at fixed low fpr) based on examples per class ($\log S$) and number of classes ($\log C$), which follows the functional form of the theoretically derived power-law. We show both very good fit on the training data as well as good prediction quality on unseen data from a different feature extractor and when fine-tuning other parameterisations (see Figure 4).
2 Background
@@
Membership inference attacks (MIAs) aim to infer whether a particular sample was part of the training set of the targeted model (Shokri et al., 2017). Thus, they can be used to determine lower bounds on the privacy leakage of models to complement the theoretical upper bounds obtained through differential privacy.
@@
Likelihood Ratio attack (LiRA; Carlini et al., 2022) While many different MIAs have been proposed (Hu et al., 2022), in this work we consider the Likelihood Ratio Attack (LiRA). LiRA is a strong attack that assumes an attacker that has black-box access to the attacked model, knows the training data distribution, the training set size, the model architecture, hyperparameters and training algorithm. Based on this information, the attacker can train so-called shadow models (Shokri et al., 2017) which imitate the model under attack but for which the attacker knows the training dataset.
@@
LiRA exploits the observation that the loss function value used to train a model is often lower for examples that were part of the training set compared to those that were not. For a target sample $(x,y_{x})$, LiRA trains the shadow models: 1 with $(x,y_{x})$ as a part of the training set ($(x,y_{x})\in\mathcal{D}$) and 2 without $x$ in the training set ($(x,y_{x})\notin\mathcal{D}$) . After training the shadow models, $(x,y_{x})$ is passed through the shadow models, and based on the losses (or predictions) two Gaussian distributions are formed: one for the losses of $(x,y_{x})\in\mathcal{D}$ shadow models, and one for the $(x,y_{x})\notin\mathcal{D}$. Finally, the attacker computes the loss for the point $x$ using the model under attack and determines using a likelihood ratio test on the distributions built from the shadow models whether it is more likely that $(x,y_{x})\in\mathcal{D}$ or $(x,y_{x})\notin\mathcal{D}$. We use an optimization by Carlini et al. (2022) for performing LiRA for multiple models and points without training a computationally infeasible number of shadow models. It relies on sampling the shadow datasets in a way that each sample is in expectation half of the time included in the training dataset of a shadow model and half of the time not. At attack time each model will be attacked once using all other models as shadow models.
Robust Membership Inference Attack (RMIA; Zarifzadeh et al., 2024) Recently Zarifzadeh et al. (2024) proposed a new MIA algorithm called RMIA, which aims to improve performance when the number of shadow models is limited. Similar to LiRA, RMIA is based on shadow model training and computing the attack statistics based on a likelihood ratio. The main difference to LiRA is that RMIA does not compute the likelihood ratio based on aggregated IN/OUT statistics, but instead compares the target data point against random samples $(z,y_{z})$ from the target data distribution. After computing the likelihood ratios over multiple $(z,y_{z})$ values, the MIA score is estimated as a proportion of the ratios exceeding a preset bound. This approach makes RMIA a more effective attack when the number of shadow models is low.